# Learning to Deliberate: Meta-policy Collaboration for Agentic LLMs with Multi-agent Reinforcement Learning

### v1

@@ Abstract
Multi-agent systems of large language models (LLMs) show promise for complex reasoning, but their effectiveness is often limited by fixed collaboration protocols. These frameworks typically focus on macro-level orchestration while overlooking agents’ internal deliberative capabilities. This critical meta-cognitive blindspot treats agents as passive executors unable to adapt their strategy based on internal cognitive states like uncertainty or confidence. We introduce the Meta-Policy Deliberation Framework (MPDF), where agents learn a decentralized policy over a set of high-level meta-cognitive actions: Persist, Refine, and Concede. To overcome the instability of traditional policy gradients in this setting, we develop SoftRankPO, a novel reinforcement learning algorithm. SoftRankPO stabilizes training by shaping advantages based on the rank of rewards mapped through smooth normal quantiles, making the learning process robust to reward variance. Experiments show that MPDF with SoftRankPO achieves a a 4–5% absolute gain in average accuracy across five mathematical and general reasoning benchmarks compared to six state-of-the-art heuristic and learning-based multi-agent reasoning algorithms. Our work presents a paradigm for learning adaptive, meta-cognitive policies for multi-agent LLM systems, shifting the focus from designing fixed protocols to learning dynamic, deliberative strategies.
@@ Introduction
As the limitations of single-agent reasoning become apparent in complex tasks (Park et al. 2023; Hong et al. 2023), the research frontier is shifting toward multi-agent systems (MAS), where multiple large language model (LLM) agents engage in collaborative reasoning to improve accuracy and robustness (Wang et al. 2023; Srivatsa, Maurya, and Kochmar 2024). Their effectiveness often depends on well-designed collaboration protocols, such as multi-turn debates and peer review mechanisms (Du et al. 2023; Liu et al. 2024). However, static protocols prevent agents from adapting their behavior to the unique demands of a given problem and to evolving context (Du et al. 2024; Yue et al. 2025).
@@
To address these deficiencies, we introduce the Meta-Policy Deliberation Framework (MPDF), which formulates multi-agent collaboration as a decentralized partially observable Markov decision process (Dec-POMDP). Unlike prior approaches relying on static heuristics or predefined rules, our framework equips each agent with a lightweight, decoupled policy network. Each agent explicitly reasons about its internal cognitive state and strategically selects from discrete deliberation acts, such as Persist, Refine, or Concede. This meta-cognitive capability enables agents to dynamically calibrate their behavior according to evolving confidence and situational context.
To robustly optimize this policy under sparse and noisy team-level feedback, we introduce SoftRankPO, a scale-resilient reinforcement learning algorithm. SoftRankPO stabilizes policy gradients by converting raw rewards into rank-based advantages derived from smooth Gaussian percentiles, effectively mitigating the reward-scale sensitivity and high-variance challenges. Our main contributions are:
@@
Recent advances in large language models (LLMs) have led to growing interest in multi-agent systems (Abdelnabi et al. 2023; Wu et al. 2024; Motwani et al. 2024; Ishibashi and Nishimura 2024; Trirat, Jeong, and Hwang 2024; Yan et al. 2025; Dai et al. 2025). Collaborative multi-agent LLM frameworks for reasoning and planning (Qian et al. 2023; Zhang et al. 2023; Zheng et al. 2023; Chen, Han, and Zhang 2024; Zhang et al. 2025a; Zu et al. 2025) mainly follow two paradigms. First, prestructured approaches rely on fixed coordination schemes (Liu et al. 2023), like debate (Du et al. 2023; Liu et al. 2024) and peer review pipelines. These systems adopt static topologies to guide communication (Du et al. 2023; Qian et al. 2024). Second, self-organizing methods adapt interactions during inference by restructuring collaboration graphs (Hu, Lu, and Clune 2024; Shang et al. 2024; Zhang et al. 2024a). Some frame workflow design as program synthesis over modules (Zhuge et al. 2024; Zhang et al. 2024b), or assign dynamic roles across agents (Hu et al. 2024; Yue et al. 2025). However, these methods mainly rely on heuristic control or static policies (Chan et al. 2023; Talebirad and Nadiri 2023; Wei, Shan, and Li 2025).
@@
Reinforcement learning (RL) for multi-agent systems aim to train coordinated, role-specialized agents beyond static prompting or supervised tuning (Chen et al. 2022a, b; Jia et al. 2025; Zhu et al. 2025; Peiyuan et al. 2024; Park et al. 2025; Liao et al. 2025; Han et al. 2025; Wan et al. 2025). These methods demonstrate that multi-agent RL can significantly improve collaboration and generalization over prompting-based systems (Jiang et al. 2025; Alsadat and Xu 2024; Lin et al. 2025). To support scalable optimization, frameworks like MARTI (Zhang et al. 2025b) and MHGPO (Chen et al. 2025a) adopt centralized coordination with decentralized policy learning. Others, such as ReSo (Zhou et al. 2025) and Verco (Li et al. 2024) introduce dynamic agent selection and interpretable communication protocols. Together, these works represent a growing trend toward training LLM-based MAS that learn to communicate and adapt under reinforcement objectives (Wang et al. 2024; Zhang et al. 2024c; Sun, Huang, and Pompili 2024; Chen et al. 2025b).
@@
We define an agent’s local state representation $z_{t}^{i}$ on a structured, low-dimensional vector summarizing its meta-cognitive status. This design choice forces the policy to learn abstract strategic reasoning rather than superficial text matching. The state $z_{t}^{i}$ is composed of three key components, with full feature specifications and prompt templates detailed in Appendix B.
Decision Schema ($z_{ans}$) represents the agent’s conclusive output, providing a concise summary of what the agent believes. It is extracted directly from the agent’s textual response by parsing the final answer into a standardized format (e.g., a scalar value for mathematical problems or a one-hot vector for multiple-choice questions).
Reasoning Profile ($z_{prof}$) is a vector of self-reported features that summarizes how the agent reached its belief. These features are produced by the agent LLM in a structured format as part of its single generative step, capturing attributes like the number of reasoning steps, the logical operators employed (e.g., algebraic, arithmetic), and a self-assessed confidence score. This co-generation process yields meta-data without incurring additional inference cost.
Introspective Confidence ($z_{conf}$) is generated via a dedicated introspection step, where a critic model, specifically the same LLM backbone acting in a evaluation capacity, is prompted to assess the correctness of the agent’s full reasoning trace given the original problem. The critic’s categorical judgment (e.g., “Correct” or “Incorrect”) is then encoded into a vector, providing a signal for further deliberation.
An agent’s full observation $o_{t}^{i}$ for decision-making is composed of its own local meta-cognitive state and those of its peers: $o_{t}^{i}=(z_{t}^{i},\{z_{t}^{j}\}_{j\neq i})$. The policy network, as a core component of our deliberation framework, then utilizes a cross-attention mechanism to integrate its introspective assessment with the social context provided by its peers. Please see Appendix for the full policy module structure.
@@
Unlike in traditional control tasks with smooth feedback, the rewards derived from deliberation are inherently sparse, high-variance, and arise from heterogeneous action outcomes. The Refine action, in particular, which depends on the stochastic generative output of an LLM, introduces significant uncertainty into the learning process. This environment yields a reward signal that is often heavy-tailed and whose scale can vary by orders of magnitude across different problem instances, compromising standard policy gradient methods, including PPO and even the more advanced GRPO. The stability of these algorithms hinges on the quality of the advantage function.In practice, rewards across tasks vary by orders of magnitude, and even simple affine rescaling $R\mapsto\alpha R+\beta$ can lead to arbitrarily inflated or collapsed gradients. The gradient variance satisfies $\operatorname{Var}[\nabla\mathcal{L}]\propto\operatorname{Var}[R]$, making learning unstable under batch-specific reward shifts. Although techniques such as baseline subtraction or batch normalization attempt to control for variance, they do not eliminate the underlying dependence on reward scale, which remains a brittle foundation in collaborative systems where agents operate with heterogeneous criteria and subjective feedback.
@@
We introduce SoftRankPO, a KL–regularised policy‑optimization method that aligns a trainable policy with ranked preferences. By divorcing learning signals from raw reward magnitudes, SoftRankPO avoids scale‑induced instability and eliminates the need for separate value critics, while still operating inside a familiar KL trust‑region.
Consider a state $s$ with reward vector $\mathbf{R}=(R_{1},\dots,R_{K})$ for the $K$ candidate actions. We first convert ordinal information into a standardised score through the mapping
$p_{i}=(\frac{\operatorname{rank}(R_{i})+\tfrac{1}{2}}{K})^{\tau},\qquad A_{i}=\frac{A_{i}^{\star}-\bar{A}^{\star}}{\mathrm{std}(\mathbf{A}^{\star})+\varepsilon},$ (2)
where $\Phi^{-1}$ denotes the inverse standard normal CDF of $\mathcal{N}(0,1)$, $A_{i}^{\star}=\Phi^{-1}(p_{i})$ and $\bar{A}^{\star}=\tfrac{1}{K}\sum_{j}A_{j}^{\star}$. The temperature parameter $\tau>0$ controls the contrast between high- and low-ranked actions, enabling smoother or sharper advantage shaping. Equation (2) leaves the relative ranking of rewards intact, removes any dependence on affine transformations of $\mathbf{R}$, and produces zero‑mean, low‑variance scores whenever $K$ is reasonably large.
Theorem 1 (Zero‑mean and bounded‑variance of SoftRank scores).
Let $\mathbf{R}=(R_{1},\dots,R_{K})$ be a reward vector with distinct entries, and let $A_{i}$ be the SoftRank scores obtained from Eq. (2) with temperature $\tau>0$. Then
[…]
Through this mechanism, we obtain a stable and reliable advantage signal. Next, we show how to incorporate this new signal into a constrained optimization objective.
KL‑constrained Rank-matching Objective
Let $\pi_{\mathrm{ref}}$ be a frozen reference policy obtained, for example, from supervised fine‑tuning. Under SoftRankPO, we maximize
@@
where $\beta>0$ controls the KL trust‑region and $\mathcal{H}$ denotes Shannon entropy. Rewriting in terms of the implicit KL reward, $R_{\theta}(s,a)=\beta\log\frac{\pi_{\theta}(a|s)}{\pi_{\mathrm{ref}}(a|s)}$. Exploiting $\sum_{i}A_{i}(s)=0$ shows that maximizing $J(\theta)$ is equivalent to minimizing the following rank-matching loss:
$\mathcal{L}_{\mathrm{SR}}(\theta)=\mathbb{E}_{s}\!\Bigl{[}\frac{1}{K}\sum_{i=1}^{K}\bigl{(}\,R_{\theta}(s,a_{i})-\bar{R}_{\theta}(s)-R(s,a_{i})+\bar{R}(s)\bigr{)}^{\!2}\Bigr{]}.$ (4)
Hence SoftRankPO learns logits whose relative shape mirrors that of the true ranked rewards, without requiring any partition function. SoftRankPO can be interpreted either as a KL‑regularised policy optimizer or as a rank‑regression procedure that aligns implicit rewards with true rewards.
Closed‑form Gradient and Structural Decomposition.
Unlike classical policy gradient methods that rely on stochastic returns or value estimators, SoftRankPO admits a fully differentiable loss function grounded in observable reward differences, enabling an explicit derivation of the gradient with respect to policy parameters while also revealing the intrinsic structure of the update rule.
Since the advantage $\mathbf{A}$ has zero mean, the partition function derivative vanishes and
$\nabla_{\theta}\mathcal{L}_{\mathrm{SR}}(\theta)=-\frac{1}{\beta}\sum_{s,i}\bigl{[}(R-\bar{R})-(R_{\theta}-\bar{R}_{\theta})\bigr{]}\,\nabla_{\theta}\log\pi_{\theta}(a_{i}|s).$ (5)
Equation (5) can be recognised as the gradient of the mean‑squared error in Eq (4). Rearranging terms yields the following triple decomposition:
$\mathcal{L}_{\mathrm{SR}}(\theta)$ $=-\mathbb{E}_{s,a}\!\bigl{[}(R-\bar{R})\log\pi_{\theta}\bigr{]}\;+\;\frac{1}{2}\,\mathbb{E}_{s}\!\bigl{[}\operatorname{Var}_{a}[\log\pi_{\theta}]\bigr{]}$
$\quad\;+\;\mathbb{E}_{s}\!\bigl{[}\operatorname{Cov}_{a}[\log\pi_{\theta},\log\pi_{\mathrm{ref}}]\bigr{]},$ (6)
which simultaneously captures advantage weighting, entropy‑variance control, and KL trust‑region coupling within a single self‑consistent loss.
Global Optimality under the KL Constraint.
The strict convexity of $\mathcal{L}_{\mathrm{SR}}$ in $R_{\theta}$ yields a unique solution.
Theorem 2 (Unique KL‑constrained optimum).
The loss $\mathcal{L}_{\mathrm{SR}}(\theta)$ admits the unique minimizer
$\pi_{\theta}^{\star}(a|s)=\frac{\pi_{\mathrm{ref}}(a|s)\exp\{\beta^{-1}R(s,a)\}}{\sum_{a^{\prime}}\pi_{\mathrm{ref}}(a^{\prime}|s)\exp\{\beta^{-1}R(s,a^{\prime})\}},$
satisfied by the KL‑constrained reward‑maximizing policy.
This result guarantees that SoftRankPO remains faithful to the same trust‑region geometry exploited by PPO and DPO, while enjoying robustness to reward scaling.
Variance control and convergence rate.
We compare single‑batch gradient variance with the Group‑Relative Policy Optimisation (GRPO) baseline and state a first‑order convergence rate under standard smoothness assumptions.
Theorem 3 (Variance Dominance and Convergence).
Let $\widehat{g}_{\mathrm{SR}}$ and $\widehat{g}_{\mathrm{GR}}$ denote stochastic gradients computed from the same mini‑batch for SoftRankPO and GRPO, respectively. If the centred reward variance satisfies $\sigma_{R}^{2}\geq 1$, then
$\operatorname{Var}\!\bigl{[}\widehat{g}_{\mathrm{SR}}\bigr{]}\leq\bigl{(}1+\tfrac{1}{K-1}\bigr{)}\,\operatorname{Var}\!\bigl{[}\widehat{g}_{\mathrm{GR}}\bigr{]}.$
Stochastic gradient descent with step sizes $\eta_{t}=\eta_{0}/\sqrt{t}$} achieves $\min_{0\leq t<T}\mathbb{E}\|\nabla_{\theta}\mathcal{L}_{\mathrm{SR}}(\theta_{t})\|^{2}=\mathcal{O}(T^{-1/2}).$
The algorithm remains efficient with $O(K\log K)$ time complexity; supporting proofs are detailed in Appendix A.
@@
The step-wise delta reward quantifies an agent’s immediate self-improvement, defined as the change in correctness of its own output before and after an action: $r_{\text{local}}=\mathbb{I}[\text{Correct}^{\text{after}}]-\mathbb{I}[\text{Correct}^{\text{before}}]$}, encouraging agents to increase their individual answer quality. The lookahead consensus reward captures an agent’s contribution to the final consensus, defined by its marginal impact on the outcome. Let $\text{Consensus}^{\text{full}}$ and $\text{Consensus}^{-i}$ denote the correctness of the final group decision with and without agent $i$’s input. The global reward is computed as $r_{\text{global}}=\mathbb{I}[\text{Consensus}^{\text{full}}]-\mathbb{I}[\text{Consensus}^{-i}]$}.
@@
Our training pipeline follows a two-stage process that integrates supervised guidance with variance-robust reinforcement learning, enabling stable convergence of the deliberation strategy. First, for robust initialization, we pre-train the policy network $\pi_{\theta}$ on an expert dataset $\mathcal{D}_{\text{sft}}=\{(s,a^{\star})\}$, where labels $a^{\star}$ are derived from oracle solutions, by minimizing a standard cross-entropy loss, yielding a strong reference policy $\pi_{\text{sft}}$. Subsequently, we generate a static offline corpus $\mathcal{D}_{\text{off}}$ by executing rollouts with this $\pi_{\text{sft}}$ policy. This corpus stores tuples of encountered meta-cognitive states and the resulting group-wise reward vectors, i.e., $\mathcal{D}_{\text{off}}=\{(s,\mathbf{R})\}$. Second, we fine-tune $\pi_{\text{sft}}$ on $\mathcal{D}_{\text{off}}$ using our SoftRankPO objective (Eq. 3), aligning the policy with the stable, rank-based advantages derived from the offline data. The KL-divergence term ensures that the policy remains within a trust region of the sensible SFT initialization.
@@
We evaluate our framework on three standard reasoning tasks for LLM agents. (i) Mathematical reasoning is assessed on GSM8K, MATH, AIME, and AMC. (ii) General reasoning is measured with the MMLU benchmark. (iii) Code generation is evaluated on HumanEval. For all data sets, we follow the original splits and report the official metrics: Solve Rate for GSM8K / MATH / AIME / AMC, Accuracy in MMLU and Pass@1 in HumanEval.
@@
We compare MPDF against a comprehensive suite of baselines, including single-agent methods: Vanilla, CoT and SC; multi-agent debate frameworks: LLM-Debate, PHP and DyLan; and dynamic workflow systems: AgentPrune, AFlow and GPTSwarm.
@@
Vanilla N Single 72.71 66.91 3.33 8.43 57.83 48.17 42.90
CoT N Single 74.00 68.18 3.33 12.05 61.60 51.83 45.17
SC N Single 80.21 73.09 4.44 12.05 68.64 56.10 49.09
PHP Y Debate 80.36 74.36 4.44 15.67 68.72 56.70 50.04
LLM-Debate Y Debate 83.47 78.91 5.56 18.07 67.41 58.54 51.99
DyLAN Y Debate 82.63 76.55 6.67 19.28 66.84 60.37 51.50
GPTSwarm Y Dynamic 84.07 73.64 5.56 15.66 69.48 57.32 51.14
AgentPrune Y Dynamic 84.53 79.45 4.44 14.46 68.82 57.92 51.60
AFlow Y Dynamic 83.40 75.63 3.33 10.84 69.07 62.16 50.74
Ours: SFT Y Deliberation 83.55 79.27 5.56 16.87 68.73 58.54 52.09
Ours: SFT + PPO Y Deliberation 82.79 79.82 4.44 13.25 68.15 56.62 50.85
Ours: SFT + GRPO Y Deliberation 85.52 81.45 7.78 20.48 70.22 60.98 54.40
Ours: SFT + SoftRankPO Y Deliberation 86.35 82.72 7.78 22.89 70.87 61.59 55.37
Table 1: Comparison of performance across reasoning tasks. “MA” denotes the use of multiple agents. “Avg” reports the average accuracy across all applicable benchmarks. Our method consistently outperforms both single-agent and multi-agent baselines.
Model LMA-8B LMA-3B Qwen-8B Qwen-3B
Vanilla 72.71 24.72 90.90 83.62
@@
Dataset LLM-Debate Ours Improvement (%)

### v2

@@ Abstract
Multi-agent systems of large language models (LLMs) show promise for complex reasoning, but their effectiveness is often limited by fixed collaboration protocols. These frameworks typically focus on macro-level orchestration while overlooking agents’ internal deliberative capabilities. This critical meta-cognitive blindspot treats agents as passive executors unable to adapt their strategy based on internal cognitive states like uncertainty or confidence. We introduce the Meta-Policy Deliberation Framework (MPDF), where agents learn a decentralized policy over a set of high-level meta-cognitive actions: Persist, Refine, and Concede. To overcome the instability of traditional policy gradients in this setting, we develop SoftRankPO, a novel reinforcement learning algorithm. SoftRankPO stabilizes training by shaping advantages based on the rank of rewards mapped through smooth normal quantiles, making the learning process robust to reward variance. Experiments show that MPDF with SoftRankPO achieves a 4-5% absolute gain in average accuracy across six mathematical and general reasoning benchmarks compared to state-of-the-art heuristic and learning-based multi-agent reasoning algorithms. Our work presents a paradigm for learning adaptive, meta-cognitive policies for multi-agent LLM systems, shifting the focus from designing fixed protocols to learning dynamic, deliberative strategies.
@@ Introduction
As the limitations of single-agent reasoning become apparent in complex tasks (Park et al. 2023; Hong et al. 2023), the research frontier is shifting toward multi-agent systems (MAS), where multiple large language model (LLM) agents engage in collaborative reasoning to improve accuracy and robustness (Chen et al. 2025d; Ping et al. 2025a; Yang et al. 2025a). Their effectiveness often depends on well-designed collaboration protocols, such as multi-turn debates and peer review mechanisms (Du et al. 2023; Liu et al. 2024). However, static protocols prevent agents from adapting their behavior to the unique demands of a given problem and to evolving context (Du et al. 2024; Yue et al. 2025).
@@
To address these deficiencies, we introduce the Meta-Policy Deliberation Framework (MPDF), which formulates multi-agent collaboration as a decentralized partially observable Markov decision process (Dec-POMDP). Unlike prior approaches relying on static heuristics or predefined rules, our framework equips each agent with a lightweight, decoupled policy network. Each agent explicitly reasons about its internal cognitive state and strategically selects from discrete deliberation acts such as Persist, Refine, or Concede. This meta-cognitive capability enables agents to dynamically calibrate their behavior according to evolving confidence and situational context.
To robustly optimize this policy under sparse and noisy team-level feedback, we introduce SoftRankPO, a scale-resilient reinforcement learning algorithm. SoftRankPO helps stabilize policy gradients by converting raw rewards into rank-based advantages derived from smooth Gaussian percentiles, thereby mitigating reward-scale sensitivity and high-variance challenges.
Our main contributions are:
@@
Beyond large language models (Chen et al. 2025c; Ping et al. 2025b; Li et al. 2025), collaborative multi-agent LLM frameworks for reasoning and planning (Motwani et al. 2024; Ishibashi and Nishimura 2024; Trirat, Jeong, and Hwang 2024; Yan et al. 2025; Dai et al. 2025; Chen, Han, and Zhang 2024; Chang et al. 2025; Zhang et al. 2025) mainly follow two paradigms. First, prestructured approaches rely on fixed coordination schemes (Liu et al. 2023), like debate (Du et al. 2023; Liu et al. 2024) and peer review pipelines. These systems adopt static topologies such as chains, trees, or graphs to guide communication (Du et al. 2023; Qian et al. 2024). Second, self-organizing methods adapt interactions during inference by restructuring collaboration graphs using Monte Carlo tree search, evolutionary updates, and pruning (Hu, Lu, and Clune 2024; Shang et al. 2024; Zhang et al. 2024a). Some frame workflow design as program synthesis over modules (Zhuge et al. 2024; Zhang et al. 2024b), or assign dynamic roles across agents (Hu et al. 2024b; Yue et al. 2025). These methods have advanced the design space of LLM-based collaboration, but rely on heuristic control or static policies (Wei, Shan, and Li 2025). We overcome this limitation by introducing closed-loop learning to adapt collaboration based on downstream utility.
@@
Reinforcement learning (RL) (Chen et al. 2022a, b; Hu et al. 2024a) for multi-agent LLM systems aim to train coordinated, role-specialized agents beyond static prompting or supervised tuning (Zheng et al. 2023; Zhu et al. 2025; Peiyuan et al. 2024; Park et al. 2025; Liao et al. 2025; Han et al. 2025; Wan et al. 2025; Zhuang et al. 2024). For example, MAPoRL (Park et al. 2025) co-trains multiple LLMs through multi-turn debates using verifier-shaped collaborative rewards, and LERO (Wei, Shan, and Li 2025) employs LLMs in an evolutionary MARL loop to generate hybrid reward and observation functions. To support scalable optimization, frameworks like Maestro (Yang et al. 2025a) and MHGPO (Chen et al. 2025a) adopt centralized coordination with decentralized policy learning. Others, such as ReSo (Zhou et al. 2025) and Verco (Li et al. 2024) introduce dynamic agent selection and interpretable communication protocols. Meanwhile, ML-Agent (Liu et al. 2025) and RAGEN (Wang et al. 2025) explore step-wise RL and long-horizon reasoning, showing that LLM agents can self-improve across complex decision trajectories. Together, these works represent a growing trend toward training LLM-based MAS that learn to communicate and adapt under reinforcement objectives (Yang et al. 2025b; Zhang et al. 2024c; Sun, Huang, and Pompili 2024; Chen et al. 2025b).
@@
We define an agent’s local state $z_{t}^{i}$ on a structured, low-dimensional vector summarizing its meta-cognitive status. This design choice forces the policy to learn abstract strategic reasoning rather than superficial text matching. The state $z_{t}^{i}$ is composed of three key components, with full feature specifications and prompt templates detailed in Appendix B.
Decision Schema ($z_{ans}$) represents the agent’s conclusive output, providing a concise summary of what the agent believes. It is extracted directly from the agent’s textual response by parsing the final output into a standardized format (e.g., a scalar value for mathematical problems or a one-hot vector for multiple-choice questions).
Reasoning Profile ($z_{prof}$) is a vector of self-reported features that summarizes how the agent reached its belief. These features are produced by the LLM agent in a structured format as part of its single generative step, capturing attributes like the number of reasoning steps, the logical operators employed (e.g., algebraic, arithmetic), and a self-assessed confidence score. This co-generation process yields meta-data without incurring additional inference cost.
Introspective Confidence ($z_{conf}$) is generated via a dedicated introspection step, where a critic model, specifically the same LLM backbone acting in a evaluation capacity, is prompted to assess the correctness of the agent’s full reasoning trace given the original problem. The critic’s judgment, which may take the form of a graded assessment or confidence-aware verdict (e.g., leaning toward “Correct” or “Incorrect” with varying strength), is then embedded into a vector representation, providing a signal for deliberation.
An agent’s full observation $o_{t}^{i}$ for decision-making is composed of its own local meta-cognitive state and those of its peers: $o_{t}^{i}=(z_{t}^{i},\{z_{t}^{j}\}_{j\neq i})$. The policy network, as a core component of our deliberation framework, then utilizes a cross-attention mechanism to integrate its introspective assessment with the social context provided by its peers. Please see Appendix B for the full policy network structure.
@@
Learning a high-level deliberation policy $\pi_{\theta}$ within our Dec-POMDP setting is challenging because the reward signal is sparse and highly non-uniform across batches. The Refine action in particular depends on stochastic LLM generations, leading to heavy-tailed returns and rounds where almost all agents make similar choices. In these cases the within-batch standard deviation can be extremely small, while the across-batch variance remains large. Moreover, the step-wise differential gains and leave-one-out consensus rewards can amplify this effect: high-consensus rounds often exhibit near-zero within-batch variance, whereas batches with more diverse or difficult candidates induce much larger variability. Standard policy gradients, including PPO and GRPO, construct advantages directly from raw reward magnitudes, so their gradient variance is sensitive to such scale and variance heterogeneity even after per-group normalization. In regimes with near-degenerate batches, per-group standard-deviation normalization can become numerically brittle.
In practice, the effective reward scale can differ substantially across tasks and batches, and even simple affine rescaling $R\mapsto\alpha R+\beta$ can significantly inflate or shrink gradient magnitudes. Since the gradient variance satisfies $\operatorname{Var}[\nabla\mathcal{L}]\propto\operatorname{Var}[R]$, learning can become unstable under batch-specific reward shifts. Although techniques such as baseline subtraction, batch normalization, or group-wise standardization help reduce variance, they do not remove the underlying dependence on reward scale, which remains a fragile foundation in collaborative systems where agents operate with heterogeneous criteria and subjective feedback.
@@
We introduce SoftRankPO, a KL-regularized policy-optimization method that aligns a trainable policy with ranked preferences. By divorcing learning signals from raw reward magnitudes, SoftRankPO reduces sensitivity to reward scale and variance while still operating inside a familiar KL trust region.
Consider a state $s$ with reward vector $\mathbf{R}=(R_{1},\dots,R_{K})$ for the $K$ candidate actions. We first convert the ordinal information into a standardized score through the mapping
$p_{i}=\Bigl(\frac{\operatorname{rank}(R_{i})+\tfrac{1}{2}}{K}\Bigr)^{\tau},\qquad A_{i}=\frac{A_{i}^{\star}-\bar{A}^{\star}}{\mathrm{std}(\mathbf{A}^{\star})+\varepsilon},$ (2)
where $\Phi^{-1}$ denotes the inverse standard normal CDF of $\mathcal{N}(0,1)$, $A_{i}^{\star}=\Phi^{-1}(p_{i})$ and $\bar{A}^{\star}=\tfrac{1}{K}\sum_{j}A_{j}^{\star}$. The temperature parameter $\tau>0$ controls the contrast between high- and low-ranked actions, enabling smoother or sharper advantage shaping. Equation (2) preserves the relative ranking of rewards, is invariant to affine transformations of $\mathbf{R}$, and produces approximately zero-mean, bounded-variance scores whenever $K$ is reasonably large. In Appendix A we formalize these properties and show that the resulting advantages are sub-Gaussian, providing a stable and well-behaved signal for policy optimization.
KL-constrained Rank-matching Objective
Let $\pi_{\mathrm{ref}}$ be a frozen reference policy obtained, for example, from supervised fine-tuning. Under SoftRankPO, we optimize the KL-regularized objective
@@
where $\beta>0$ controls the KL trust region and $\mathcal{H}$ denotes Shannon entropy. Writing this in terms of the implicit KL-induced reward $R_{\theta}(s,a)=\beta\log\frac{\pi_{\theta}(a\mid s)}{\pi_{\mathrm{ref}}(a\mid s)}$ and using the zero-mean property $\sum_{i}A_{i}(s)=0$, we show in Appendix A that maximizing $J(\theta)$ is equivalent (up to constants) to minimizing the following rank-matching loss:
$\mathcal{L}_{\mathrm{SR}}(\theta)=\mathbb{E}_{s}\!\Bigl[\frac{1}{K}\sum_{i=1}^{K}\bigl(\,R_{\theta}(s,a_{i})-\bar{R}_{\theta}(s)-R(s,a_{i})+\bar{R}(s)\bigr)^{\!2}\Bigr],$ (4)
where $\bar{R}(\cdot)$ and $\bar{R}_{\theta}(\cdot)$ denote within-state averages. Hence SoftRankPO learns logits whose relative shape mirrors that of the true ranked rewards, without requiring explicit partition functions. It can be interpreted either as a KL-regularized policy optimizer or as a rank-regression procedure that aligns implicit rewards with task rewards.
Optimization behaviour and theoretical properties.
The SoftRank advantages are centered and have bounded variance, which implies that the partition-function term cancels in the gradient and the update admits a clean analytical form. In Appendix A we derive the gradient of $\mathcal{L}_{\mathrm{SR}}(\theta)$, show how it decomposes into advantage-weighted learning, entropy-variance control, and KL trust-region coupling, and restate a standard result characterizing the KL-constrained optimum. We also provide a variance comparison with GRPO and a first-order convergence rate under common smoothness assumptions.
@@
The step-wise delta reward quantifies an agent’s immediate self-improvement, defined as the change in correctness of its own output before and after an action: $r_{\text{local}}=\mathbb{I}[\text{Correct}^{\text{after}}]-\mathbb{I}[\text{Correct}^{\text{before}}]$, encouraging agents to increase their individual answer quality. This step-wise shaping is conceptually related to prior MARL work on social deduction games, where agents are rewarded for improving others’ answers (Sarkar et al. 2025), but here we focus on self-improvement and combine it with a consensus-based marginal term.
The lookahead consensus reward captures an agent’s contribution to the final consensus, defined by its marginal impact on the outcome. Let $\text{Consensus}^{\text{full}}$ and $\text{Consensus}^{-i}$ denote the correctness of the final group decision with and without agent $i$’s input. The global reward is computed as $r_{\text{global}}=\mathbb{I}[\text{Consensus}^{\text{full}}]-\mathbb{I}[\text{Consensus}^{-i}]$, and in practice $\text{Consensus}^{-i}$ is recomputed from the cached responses in the same round without requiring additional rollouts.
@@
Our training pipeline follows a two-stage process that integrates supervised guidance with variance-robust reinforcement learning, enabling stable convergence of the deliberation strategy. First, for robust initialization, we pre-train the policy network $\pi_{\theta}$ on an expert dataset $\mathcal{D}_{\text{sft}}=\{(s,a^{\star})\}$, where labels $a^{\star}$ are derived from oracle solutions, by minimizing a standard cross-entropy loss. This yields a supervised policy $\pi_{\text{sft}}$, which we also use as the fixed reference policy in the KL regularization, i.e., $\pi_{\text{ref}}=\pi_{\text{sft}}$ in Eq. 3. We initialize $\pi_{\theta}$ from $\pi_{\text{sft}}$ before reinforcement learning. Subsequently, we generate a static offline corpus $\mathcal{D}_{\text{off}}$ by executing rollouts with $\pi_{\text{sft}}$. This corpus stores tuples of encountered meta-cognitive states and the resulting group-wise reward vectors, i.e., $\mathcal{D}_{\text{off}}=\{(s,\mathbf{R})\}$. In the second stage, we fine-tune $\pi_{\theta}$ on $\mathcal{D}_{\text{off}}$ using our SoftRankPO objective (Eq. 3), aligning the policy with the stable, rank-based advantages derived from the offline data. The KL-divergence term ensures that the updated policy remains within a trust region of the sensible SFT initialization.
@@
We evaluate our framework on three standard reasoning tasks for LLM agents (Hendrycks et al. 2020; Cobbe et al. 2021; Hendrycks et al. 2021; Chen 2021; Du et al. 2023). (i) Mathematical reasoning is assessed on GSM8K, MATH, AIME, and AMC. (ii) General reasoning is measured with the MMLU benchmark. (iii) Code generation is evaluated on HumanEval. For all data sets, we follow the original splits and report the official metrics: Solve Rate for math datasets, Accuracy in MMLU and Pass@1 in HumanEval.
@@
We compare MPDF against a comprehensive suite of baselines, including single-agent methods: Vanilla, CoT and SC; multi-agent sequential and debate frameworks: PHP and LLM-Debate; and dynamic workflow systems: DyLan, AgentPrune, AFlow and GPTSwarm.
@@
Vanilla Single 72.71 66.91 3.33 8.43 57.83 48.17 42.90
CoT Single 74.00 68.18 3.33 12.05 61.60 51.83 45.17
SC Single 80.21 73.09 4.44 12.05 68.64 56.10 49.09
PHP Sequential 80.36 74.36 4.44 15.67 68.72 56.70 50.04
LLM-Debate Debate 83.47 78.91 5.56 18.07 67.41 58.54 51.99
DyLAN Dynamic 82.63 76.55 6.67 19.28 66.84 60.37 51.50
GPTSwarm Dynamic 84.07 73.64 5.56 15.66 69.48 57.32 51.14
AgentPrune Dynamic 84.53 79.45 4.44 14.46 68.82 57.92 51.60
AFlow Dynamic 83.40 75.63 3.33 10.84 69.07 62.16 50.74
MPDF: SFT Deliberation 83.55 79.27 5.56 16.87 68.73 58.54 52.09
MPDF: SFT + PPO Deliberation 82.79 79.82 4.44 13.25 68.15 56.62 50.85
MPDF: SFT + GRPO Deliberation 85.52 81.45 7.78 20.48 70.22 60.98 54.40
MPDF: SFT + SoftRankPO Deliberation 86.35 82.72 7.78 22.89 70.87 63.82 55.74
Table 1: Comparison of performance across reasoning tasks. “MA” denotes the use of multiple agents (✓) versus single-agent baselines (✗). “Avg” reports the average accuracy across all applicable benchmarks. Our method consistently outperforms both single-agent and multi-agent baselines.
Model LMA-8B LMA-3B Qwen-7B Qwen-3B
Vanilla 72.71 38.64 90.90 83.62
@@
Dataset LLM-Debate MPDF Improvement (%)